### v1

# Leveraging Label Information for Stealthy Data Stealing in Vertical Federated Learning

Abstract.
Vertical Federated Learning (VFL) is a collaborative learning paradigm designed for scenarios where multiple clients share disjoint features of the same set of data samples. Albeit a wide range of applications, VFL is faced with privacy leakage from data reconstruction attacks. These attacks generally fall into two categories: honest-but-curious (HBC), where adversaries steal data while adhering to the protocol; and malicious attacks, where adversaries breach the training protocol for significant data leakage. While most research has focused on HBC scenarios, the exploration of malicious attacks remains limited.
Launching effective malicious attacks in VFL presents unique challenges: 1) Firstly, given the distributed nature of clients’ data features and models, each client rigorously guards its privacy and prohibits direct querying, complicating any attempts to steal data; 2) Existing malicious attacks alter the underlying VFL training task, and are hence easily detected by comparing the received gradients with the ones received in honest training. To overcome these challenges, we develop DMAVFL, a novel attack strategy that evades current detection mechanisms. The key idea is to integrate a discriminator with auxiliary classifier that takes a full advantage of the label information (which was completely ignored in previous attacks): on one hand, label information helps to better characterize embeddings of samples from distinct classes, yielding an improved reconstruction performance; on the other hand, computing malicious gradients with label information better mimics the honest training, making the malicious gradients indistinguishable from the honest ones, and the attack much more stealthy. Our comprehensive experiments *demonstrate that DMAVFL significantly outperforms existing attacks, and successfully circumvents SOTA defenses for malicious attacks. Additional ablation studies and evaluations on other defenses further underscore the robustness and effectiveness of DMAVFL.
ccs: Security and privacy Distributed systems securityccs: Computing methodologies Machine learning
1. Intoduction
Federated learning (FL) is an emerging privacy-preserving collaborative learning paradigm, which allows multiple distributed clients to securely train a model without sharing private data (Yang et al., 2019b). FL can be categorized into horizontal FL (HFL) and vertical FL (VFL) (Liu et al., 2024). In HFL, different clients possess distinct sets of training samples, yet all samples share the same feature space. In contrast, VFL is applicable to scenarios where clients have the same sample identifiers but possess disjoint feature spaces, e.g., an finance company may hold investment records for a set of customers, while a bank may have the expenditure details for the same set of customers. Due to its capability to integrate diverse data sources, VFL is increasingly gaining attention, and has been applied in various fields including finance (Chen et al., 2020), healthcare (Poirot et al., 2019), and recommendation systems (Yang et al., 2020).
In VFL (Liu et al., 2022; Chen et al., 2020; Li et al., 2023; Thapa et al., 2022; Feng and Yu, 2020; Wu et al., 2020; Yang et al., 2019a), an active client holds both labels and partial features, while multiple passive clients hold disjoint features for the same samples. VFL typically employs a split learning framework (Vepakomma et al., 2018; Abuadbba et al., 2020; Thapa et al., 2022; Kim et al., 2020), and utilizes separate models to process these disjoint features. Specifically, a split VFL system comprises one top model and several bottom models: each passive client operates a bottom model, whereas the active client manages both a bottom model and a top model (see Figure 1). During the training phase, for each batch of training samples, all passive and active clients compute embeddings using their respective bottom models; these embeddings are then transmitted to the active client, who outputs the predictions with the top model and computes the loss from the labels. Following this, the gradients for embeddings are sent back to all clients to update their bottom models.
Figure 1. Illustration of a VFL system with one active client and two passive clients. The area enclosed by the red dashed lines contains the information accessible to and the actions taken by the active client. In a data reconstruction attack, the malicious active client intends to recover the private features of the target passive clients.
While VFL emphasizes privacy by sharing embeddings instead of raw data, emerging data reconstruction attacks have revealed potential vulnerabilities that could lead to privacy leakage within VFL models. Such attacks are categorized into honest-but-curious (HBC) and malicious attacks. HBC adversaries stealthily extract private features while adhering to the protocol, whereas malicious adversaries actively manipulate the protocol to steal private features. Notable methods like GRNA (Luo et al., 2021) and GIA (Jiang et al., 2022) focus on HBC scenarios. In GRNA, under the assumption that the adversary can query the passive clients’ models, target features could be inferred by training a generative regression network to learn correlations between target features and the adversary’s partial features, thus enabling generation of target features. GIA (Jiang et al., 2022), presupposing the adversary can inject adversarial features to the target clients, employs optimization techniques to reconstruct the target features. A malicious data reconstruction attack in VFL, called AGN, is developed in (Vu et al., 2023), where the adversary transmits malicious gradients to passive clients to induce more significant feature leakage.
A class of closely related attacks have also been developed for split learning (SL), which can be viewed as a special case of VFL. In SL, there is a clear division of roles where an active client, holding a top model and labels, but without partial features, collaborates with a single passive client who manages a bottom model and all data samples with complete features. In HBC scenario, attacks like PCAT (Gao and Zhang, 2023) and SDAR (Zhu et al., 2023) use the top model to train a shadow model on an auxiliary dataset, which is subsequently used to train a decoder for data reconstruction. In the malicious scenario, FSHA (Pasquini et al., 2021) manipulates the training objective, generating malicious gradients to achieve more accurate data reconstruction.
While malicious attacks could be more detrimental to privacy leakage in VFL, research works on malicious attacks remains limited. We identify two primary challenges to launch effective malicious attacks: 1) Distributed features and limited model access. The heterogeneous nature of feature sources across distributed clients, together with the lack of access to passive clients’ local models, makes it difficult for the adversary to infer target features using its own partial features; 2) Powerful defenses. Existing malicious attacks, such as FSHA and AGN, replace the honest training task with a malicious task, causing the gradients returned to passive clients to appear different from those in honest training, which renders these attacks easily detectable. Defensive strategies in SL, such as SplitGuard (Erdogan et al., 2022) and Gradient Scrutinizer (Fu et al., 2023), monitor incoming gradients to identify potential attacks. They are readily adapted to the VFL scenario to significantly diminish the success of malicious data reconstruction attacks.
In response to these challenges, we investigate whether a malicious attacker in VFL can circumvent defensive measures to maintain effective data reconstruction. We answer this question affirmatively, by developing a novel defense-penetrating malicious data reconstruction attack on VFL (referred to as DMAVFL). DMAVFL begins with the adversary pretraining an encoder, a decoder, and its bottom model to achieve high reconstruction performance on an auxiliary data. Following this, the adversary freezes the encoder and bottom model, and trains a discriminator with auxiliary classifier (DAC) (Hou et al., 2022) that integrates labels information to generate malicious gradients transmitted to passive clients. These malicious gradients can transfer embeddings distribution from the encoder to passive clients’ models. In data reconstruction phase, the trained decoder is leveraged to reconstruct target features from the target clients, as these passive clients’ models are meticulously guided to mimic the encoder by DAC.
To address challenge 1, DMAVFL captures correlations between the adversary’s partial features and the target features without direct queries or data injections, using DAC to generate malicious gradients that indirectly guide the passive clients’ model to leak feature information. To handle challenge 2, we use DAC to incorporate label information into the computation of the malicious loss function, generating malicious gradients almost indistinguishable from the gradients received in honest training, which helps DMAVFL to effectively penetrate current defenses. Additionally, these label information helps to better transfer the embedding distribution from the encoder to the passive clients’ models compared with a traditional discriminator, and provide extra information for achieving a smaller reconstruction error. As a result, over extensive empirical evaluations, DMAVFL is shown to achieve better reconstruction performance than SOTA attacks, and successfully circumvents all existing defenses against malicious attacks. Further ablation studies and evaluations against other defenses underscore the robustness and effectiveness of DMAVFL.
In summary, this paper makes the following main contributions:
  1. (1)
    We introduce DMAVFL, a novel defense-penetrating malicious data reconstruction attack in VFL, designed to breach existing security measures and effectively steal private features;
  2. (2)
    Unlike traditional approaches that utilize a discriminator, DMAVFL incorporates training a DAC, which helps to 1) significantly enhance the effectiveness in embedding distribution transfer and attack performance; and 2) make the malicious gradients indistinguishable from the ones with honest training;
  3. (3)
    We rigorously evaluate the attack performance of DMAVFL over four representative datasets, demonstrating that it circumvents all current defenses against malicious data reconstruction, and outperforms all SOTA attacks on VFL (and the ones adapted from SL), with and without defenses.
2. Background and Related works
In this section, we review SOTA data reconstruction attacks on VFL, and existing defense mechanisms for these attacks.
2.1. Data reconstruction attacks on VFL
Although VFL safeguards feature privacy by exchanging embeddings and intermediate gradients instead of raw features, recent research has revealed its vulnerability to data reconstruction attacks that can lead to private feature leakage. In HBC setting, the attacks described in (Luo et al., 2021) and (Yang et al., 2023) assume that an HBC adversary can query the passive clients’ bottom model. During the inference phase, the adversary covertly trains a generative model that is fed random vectors and partial data features available to the adversary, thereby producing synthetic inputs. By minimizing the discrepancy of the bottom models’ output between synthetic inputs and target features, the generative model is trained to approximate these features accurately. Notably, while the adversary in (Luo et al., 2021) can access the gradients of the target bottom model directly, the adversary in (Yang et al., 2023) employs zeroth-order optimization to estimate these gradients, thereby bypassing the need for direct gradient access.
The approach in (Jiang et al., 2022), GIA, operates under a black-box scenario where the adversarial active client lacks direct access to the passive clients’ bottom model. Despite this limitation, the adversary can strategically insert a batch of auxiliary data features into the datasets of the passive clients and subsequently collect the resulting output embeddings. The adversary then trains a shadow model using the collected auxiliary features and embeddings pairs. After training, the adversary optimizes noise to closely approximate the target features by querying this shadow model. It is essential to emphasize that these attack methodologies, while effective in reconstructing private data, depend on the adversary having either direct access to the passive clients’ models or the capability to inject data into these clients.
In malicious setting, the attack method, AGN (Vu et al., 2023), utilizes a generator to create fake target features from passive clients’ embeddings. These fake features are then evaluated by a discriminator, which distinguishes between real and fake features to generate malicious gradients. These gradients are sent back to the passive clients, aiming to train embeddings that facilitate the reconstruction of target features. Despite the innovative approach of AGN, our evaluations in the section of experiments reveal that the data reconstructed by this method is relatively coarse and remains susceptible to existing defense mechanisms.
Data reconstruction attacks on split learning. A parallel strand of privacy attacks on SL is adaptable to VFL with certain modifications. Unlike VFL which involves multiple clients, SL typically includes only one client and a server managing the bottom and top models, respectively, which can be viewed as a special case of VFL. In SL, adversarial servers are categorized as either honest-but-curious (HBC) or malicious.
For HBC settings, He et al. pioneer privacy attacks on SL in (He et al., 2019), exploring various levels of adversarial knowledge. In a black-box scenario, where the adversary lacks the knowledge of the model but can make queries, an inverse model is used to reconstruct private features. In scenarios without query capabilities, the adversary utilizes a shadow model trained with auxiliary data under the guidance of the top model. Similar with GIA, the shadow model is then used to reconstruct target features through optimization. Unsplit (Erdoğan et al., 2022) presents a scenario where the adversary knows the bottom model’s structure but lacks auxiliary data. This method involves optimizing synthetic features and the shadow model concurrently to reduce disparities between real and synthetic features’ embeddings. PCAT (Gao and Zhang, 2023) depicts an adversary with access to a batch of auxiliary data but unaware of the bottom model’s structure and parameters. The strategy involves covertly training a shadow model with the auxiliary data, guided by the top model in each training round. Then an inverse model is trained on the shadow model for data reconstruction. SDAR (Zhu et al., 2023), assuming knowledge of the bottom model’s architecture, utilizes auxiliary data and adversarial regularization to train both a shadow model and its inverse model. SDAR employs two discriminators for adversarial regularization: one differentiates between the embeddings of synthetic and genuine features, while the other identifies the data’s origin, whether it comes from the inverse model or the auxiliary data. For malicious setting, the attack method, FSHA (Pasquini et al., 2021), replaces the top model with a discriminator for aligning the bottom model with a local encoder. Subsequently, an decoder trained on the encoder can directly apply to reconstruct target features.
2.2. Defenses against data reconstruction attacks
As data leakage risks in VFL escalate, a variety of defensive strategies have been developed to defend data reconstruction attacks. These defenses, particularly against an HBC adversary, can be broadly classified into cryptographic methods, perturbation-based methods, and adversarial training methods. Cryptographic strategies include homomorphic encryption (HE)(Xu et al., 2021; Gong et al., 2023; Fu et al., 2022), secret sharing (SS)(Li et al., 2023; Huang et al., 2023), differential privacy (DP)(Wang et al., 2020; Chen et al., 2020), and hashing(Qiu et al., 2022). Employed to protect the privacy of individual embeddings in VFL, these methods, while effective, often lead to significant computation and communication overheads. Perturbation methods involve adding noise to embeddings to disturb potential reconstruction attacks, as detailed in (He et al., 2020). Another variation treats noise as trainable parameters, aiming at minimizing privacy leakage without significant degradation of model performance, as explored in (Mireshghallah et al., 2021). While these methods can mitigate data reconstruction attacks, they might reduce model accuracy. Adversarial training techniques, as discussed in (Sun et al., 2021; Vepakomma et al., 2020), train models in such a way that embeddings are less likely to leak raw features information. Although these methods are effective at reducing features leakage, they increase the computational burden and may affect the pefromance of the intended learning tasks.
As malicious attacks such as FSHA and AGN substitute the intended training task, causing significant private features leakage, recent advancements in defense methods, notably SplitGuard (Erdogan et al., 2022) and Gradient Scrutinizer (Fu et al., 2023), have been developed to specifically address these types of attacks. SplitGuard operates by requiring passive clients to intermittently replace labels in a batch with randomized labels, referred to as fake batches, during training. The idea behind SplitGuard is that if the top model is genuinely engaged in learning the intended task, the angles and distances between the gradients from fake and regular batches will differ significantly from those between two regular batches. Consequently, passive clients can detect anomalies by analyzing these differences in gradients. Unlike SplitGuard, Gradient Scrutinizer does not necessitate changes to labels. It operates by collecting and analyzing received batch gradients with identical and differing labels during each training round. The key finding is that during honest training, the gradient distances for embeddings with the different label should be greater than those for embeddings with same labels. By comparing these gradient distances, Gradient Scrutinizer can effectively detect deviations from honest training patterns, indicating potential attacks.
3. Problem Description
3.1. VFL protocol
In a VFL system, $N$ passive clients collaborate with an active client to train and utilize a VFL model. Each passive client, denoted as $n\in[N]$, operates a bottom model $f_{n}(\cdot)$. The active client manages a top model $f_{0}(\cdot)$ and a bottom model $f_{a}(\cdot)$. The training data $D_{train}:=(\boldsymbol{x}_{j})_{j=1}^{M}$ is vertically partitioned among $N+1$ clients. For each sample $j\in[M]$, the active client holds both partial features $\boldsymbol{x}_{j,a}$ and the label $y_{j}$; while each passive client $n$, $n\in[N]$, possesses a disjoint portion of the features $\boldsymbol{x}_{j,n}$.
In the training phase, for a selected batch $\mathcal{B}$, each passive client $n$ computes the corresponding embeddings $\boldsymbol{h}_{j,n}=f_{n}(\boldsymbol{x}_{j,n}),j\in\mathcal{B}$, and sends them to the active client. The active client processes its partial features to produce embeddings $\boldsymbol{h}_{j,a}=f_{a}(\boldsymbol{x}_{j,a}),j\in\mathcal{B}$. Upon concatenating all batch embeddings, $\boldsymbol{h}_{j}=[\boldsymbol{h}_{j,a},\boldsymbol{h}_{j,1},\ldots,\boldsymbol{h}_{j,N}]$, the active client completes forward propagation through the top model. Loss is computed as $L=\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}l(f_{0}(\boldsymbol{h}_{j}),y_{j})$, for some loss function (e.g., cross entropy) $l(\boldsymbol{a},b)$ between prediction $\boldsymbol{a}$ and label $b$. Gradients on the embeddings are sent back to each client for updating their respective bottom models. This process iterates until all models converge.
In the inference phase, when the prediction on a test sample is requested, each client computes an embedding using its bottom model and sends it to the active client. The active client aggregates these embeddings and uses the top model to predict the label. This approach ensures that data privacy is maintained by sharing only the embeddings rather than the raw features, and also enhances computational efficiency by leveraging the split learning mechanism.
3.2. Threat model
Capability and knowledge. The active client acts as a malicious adversary with the capability to deviate from the standard VFL training protocol. Specifically, this adversary can alter the training task and transmits malicious gradients to other clients. The adversary controls a top model $f_{0}(\cdot)$, and accesses partial data features $\boldsymbol{x}_{j,a},j\in[M]$, along with a bottom model $f_{a}(\cdot)$. Additionally, the adversary possesses an auxiliary dataset, $D_{aux}:=(\boldsymbol{x}_{i},y_{i})_{i=1}^{M_{aux}}$, which shares a similar distribution with the training data but remains distinct from it, i.e., $D_{aux}\cap D_{train}=\emptyset$. Each sample in the auxiliary dataset contains the complete set of features and the corresponding label. This assumption is congruent with previous attacks against SL (Gao and Zhang, 2023; Zhu et al., 2023; Pasquini et al., 2021) and presents a more realistic capability compared to those models that assume direct querying of passive clients’ bottom models. We explicitly do not assume any capability for the adversary to query or directly access the knowledge of the passive clients’ models (see Figure 1).
Goal and metric. The primary objective of the adversary is to reconstruct target clients’ data features, denoted as $\boldsymbol{X}_{t}=(\boldsymbol{x}_{j,n})_{n\in\mathcal{T}_{N},j\in\mathcal{T}_{X}}$, where $\mathcal{T}_{N}$ is the set of target clients with size $|\mathcal{T}_{N}|\leq N$, and $\mathcal{T}_{X}$ is the set of target samples. We denote the reconstructed features as $\tilde{\boldsymbol{X}}_{t}$, and use the mean squared error (MSE) between $\boldsymbol{X}_{t}$ and $\tilde{\boldsymbol{X}}_{t}$ over target samples as the metric to measure the effectiveness of the attack.
4. Challenges and limitations
We point out the main challenges and the limitations of the current attack methods, to achieve effective and stealthy data reconstruction under practical assumptions on VFL settings.
Challenge 1: distributed features and limited model access. VFL involves multiple clients each holding a unique subset of data features and a bottom model. This dispersed and heterogeneous environment raises challenges for the adversary, to use its own partial features to effectively reconstruct target features from various sources. Additionally, given the distributed nature of bottom models across multiple clients, each guarding its data privacy, any attempt at querying or data injection would trigger an alarm for potential privacy breach and thus, is typically prohibited. This inaccessibility of passive clients’ models complicates the adversary’s strategy to manipulate them for data leakage.
Challenge 2: powerful defenses. Malicious data reconstruction attacks are prone to detection by sophisticated defense mechanisms such as SplitGuard and Gradient Scrutinizer. One major reason is that current malicious attacks, e.g. AGN and FSHA, alter the honest training task and thus affecting the transmitted gradients. While such defenses are highly sensitive to deviations in training behavior, as they typically monitor gradient flows for anomalies that could indicate malicious activities. To make the attack more stealthy, a natural solution is to sum up the honest training task (e.g., classification) loss and the adversarial loss. To evaluate the efficacy of this modification, we employed SplitGuard as a defense mechanism to test against malicious attack strategies AGN and FSHA on MNIST dataset. As depicted in Figure 2, the detection scores for both attacks, even with the modified loss, remained below the threshold of 0.9, yielding successful detection. This outcome suggests that simply combining loss functions is insufficient to penetrate current defenses.
Figure 2. Comparison of Splitguard scores for two malicious attacks, and their variants with modified loss functions.
Limitations: 1) Existing HBC data reconstruction attacks on VFL operates under strong, often unrealistic assumptions about adversary’s capabilities, such as the ability to query the target model or inject data directly into target clients, which are impractical in real-world scenarios; 2) Furthermore, the attacks in SL focus on reconstructing features from a single target client, and do not use any partial features of the data, limiting their applicability in VFL; 3) Malicious attacks like FSHA and AGN completely ignore the label information in their designs, which is the root cause of their detectability, and incorporating the labels into attack designs can potentially further improve the reconstruction performance.
The above challenges and limitations steer us toward a compelling research question: ”Under a practical adversary model, can we develop a malicious data reconstruction attack in VFL that successfully circumvents current defense mechanisms while achieving superior reconstruction performance?” We affirmatively answer this question by developing a novel attack strategy DMAVFL, which addresses all highlighted challenges and limitations.
5. Methodology of DMAVFL
In this section, we introduce our DMAVFL strategy. As illustrated in Figure 3, DMAVFL consists of the following three steps.
  1. (1)
    Pretraining: Before initiating VFL training, the adversary sets up three local models consisting of an encoder, a decoder, and a Discriminator with Auxiliary Classifier (DAC). The pretraining step involves the concurrent training of these models. This step is critical for minimizing reconstruction loss and effectively preparing the models for the subsequent steps of the attack.
  2. (2)
    Malicious gradient generation: With the onset of VFL training, the adversary freezes the encoder to preserve its learned embeddings and replaces the conventional top model with the DAC. The DAC is instrumental in transferring the embedding distribution from the encoder into the target model and integrating label information through classification processes. This integration ensures that the malicious training is indistinguishable from honest training, enhancing the stealthiness of the attack. Utilizing the malicious loss computed through DAC, the adversary meticulously crafts malicious gradients. These gradients are then dispatched to the target clients to manipulate their models, mimicking the embedding distribution of the encoder.
  3. (3)
    Data reconstruction: In the data reconstruction phase, the adversary leverages the decoder to reconstruct private features of target clients, from embeddings uploaded by passive clients and adversary’s local embedding.
Figure 3. Workflow of DMAVFL. The opaque rectangle indicates that the model is being trained, while the transparent rectangle represents that the model is frozen and only utilized in forward propagation.
5.1. Setup
We first denote a target model $f_{p}(\cdot):=(f_{n}(\cdot))_{n\in[N]}$ as the collection of the bottom models of all passive clients. In the setup phase, the adversary prepares the following three networks in replacement of the original top model:
  1. (1)
    Encoder ($f_{e}(\cdot)$): The encoder $f_{e}(\cdot)$ is designed to mimic the functionality of the target model $f_{p}(\cdot)$. The dimensions of the encoder are carefully aligned with those of $f_{p}(\cdot)$, such that its input dimension equals to the total number of features on the passive clients, and its output dimension equals to the sum of output dimensions of all passive clients’ the bottom models.
  2. (2)
    Decoder ($f_{d}(\cdot)$): The decoder is tailored to reconstruct the target clients’ data features. It accepts concatenated embeddings as input and outputs data features of the target clients.
  3. (3)
    Discriminator with Auxiliary Classifier (DAC) ($D(\cdot)$) : The DAC is engineered to classify and discriminate between real labels $y^{+}$ and fake labels $y^{-}$, aiding in the validation of the embeddings’ authenticity and malicious gradient generation.
Additionally, the adversary maintains a bottom model $f_{a}(\cdot)$. This model processes the partial data features available to the adversary, capturing their correlations with the target features.
5.2. Pretraining
Prior to commencing VFL training, the adversary constructs an encoder-decoder structure consisting of the adversary’s bottom model $f_{a}(\cdot)$, the encoder $f_{e}(\cdot)$, and the decoder $f_{d}(\cdot)$. This setup is pretrained on the auxiliary dataset $D_{aux}$. The objective of this pretraining phase is threefold: to train the decoder to reconstruct target features, to enable the adversary’s bottom model to learn the relationship between target and adversary’s partial features, and to allow the encoder to capture the representation of private features for effective reconstruction.
In each iteration of training, the adversary selects a batch of auxiliary data, denoted as $\mathcal{B}_{aux}$. This batch is partitioned into two subsets: the adversary’s features $\boldsymbol{x}_{i,a}$ and the passive features $\boldsymbol{x}_{i,p},i\in\mathcal{B}_{aux}$, which are fed into $f_{a}(\cdot)$ and $f_{e}(\cdot)$, respectively. The resulting embeddings, $\boldsymbol{h}_{i,a}=f_{a}(\boldsymbol{x}_{i,a})$ and $\boldsymbol{h}_{i,p}=f_{e}(\boldsymbol{x}_{i,p}),i\in\mathcal{B}_{aux}$ are concatenated to form $\boldsymbol{h}_{i}=(\boldsymbol{h}_{i,a}\|\boldsymbol{h}_{i,p})$. These concatenated embeddings are then forwarded to the decoder, which produces synthetic target features $\tilde{\boldsymbol{x}}_{i,t}=f_{d}(\boldsymbol{h}_{i})$. The primary objective of the adversary is to minimize the reconstruction loss, which is computed as follows:
(1) $\mathcal{L}_{R}=\frac{1}{|\mathcal{B}_{aux}|}\sum_{i\in\mathcal{B}_{aux}}MSE(\tilde{\boldsymbol{x}}_{i,t},\boldsymbol{x}_{i,t}).$
Data: Auxiliary dataset $D_{aux}=\{\boldsymbol{x}_{i,a},\boldsymbol{x}_{i,p}\}_{i=1}^{M_{aux}}$.
Initialization: Encoder $f_{e}(\cdot)$, decoder $f_{d}(\cdot)$, adversary’s bottom model $f_{a}(\cdot)$.
while $f_{e}(\cdot)$, $f_{d}(\cdot)$, and $f_{a}(\cdot)$ not converge do
      The adversary select a batch $\mathcal{B}_{aux}$ from $D_{aux}$;
      Compute $\boldsymbol{h}_{i,a}=f_{a}(\boldsymbol{x}_{i,a}),i\in\mathcal{B}_{aux}$, and $\boldsymbol{h}_{i,p}=f_{e}(\boldsymbol{x}_{i,p}),i\in\mathcal{B}_{aux}$;
      Compute $\tilde{\boldsymbol{x}}_{i,t}=f_{d}([\boldsymbol{h}_{i,a}\|\boldsymbol{h}_{i,p}]),i\in\mathcal{B}_{aux}$;
      $\mathcal{L}_{R}\leftarrow\frac{1}{|\mathcal{B}_{aux}|}\sum_{i\in\mathcal{B}_{aux}}MSE(\tilde{\boldsymbol{x}}_{i,t},\boldsymbol{x}_{i,t})$;
      $\boldsymbol{\nabla}_{e}\leftarrow$ Gradient($\mathcal{L}_{R},f_{e}(\cdot)$);
      $\boldsymbol{\nabla}_{a}\leftarrow$ Gradient($\mathcal{L}_{R},f_{a}(\cdot)$);
      $\boldsymbol{\nabla}_{d}\leftarrow$ Gradient($\mathcal{L}_{R},f_{d}(\cdot)$);
      $f_{e}(\cdot)\leftarrow\text{Model\_update}(f_{e}(\cdot),\boldsymbol{\nabla}_{e}$);
      $f_{a}(\cdot)\leftarrow\text{Model\_update}(f_{a}(\cdot),\boldsymbol{\nabla}_{a}$);
      $f_{d}(\cdot)\leftarrow\text{Model\_update}(f_{d}(\cdot),\boldsymbol{\nabla}_{d}$);
end while
Algorithm 1 Pretraining.
The pretraining process is detailed in Algorithm 1, where the function Grdient($L,\boldsymbol{\theta}$) computes the gradients of the loss $L$ with respect to parameters $\boldsymbol{\theta}$; and the function Model_update($f(\cdot),\nabla_{\boldsymbol{\theta}}$) represents the operation of updating the model $f(\cdot)$ using gradients $\nabla_{\boldsymbol{\theta}}$.
5.3. Malicious gradient generation
After the pretraining, the adversary intends to transfer the embeddings distribution from the encoder $f_{e}(\cdot)$ to the target model $f_{p}(\cdot)$. Since the adversary lacks direct access to modify $f_{p}(\cdot)$, they resort to transmitting malicious gradients to the passive clients, which are designed to guide $f_{p}(\cdot)$ to mimic the behavior of $f_{e}(\cdot)$. In this sequel, for ease of exposition, we describe all operations with $f_{p}(\cdot)$, which means that the same operation is simultaneously performed on the bottom models of all passive clients.
An intuitive solution involves employing a discriminator to distinguish between embeddings from the encoder (labeled as real) and those produced by the target model (labeled as fake). The target model is then trained to deceive the discriminator into recognizing its embeddings as real, thus aligning its distribution with that of the encoder. This strategy focuses on minimizing the Jensen–Shannon divergence (JS) (Fuglede and Topsoe, 2004) between the distributions of the embeddings from the auxiliary and training data, i.e., $\textup{JS}(P_{\boldsymbol{H}_{aux}}\|P_{\boldsymbol{H}_{train}})$, which facilitates the transfer of embeddings distribution.
However, a primary concern is that the discriminator approach completely ignores the label information available at the adversary, which may lead to 1) sub-optimal reconstruction performance; and 2) detection by defenses that can tell the difference between gradients computed using and without using labels. To address these issues, we employ DAC, $D(\cdot)$, defined in (2), which not only differentiates between real and fake embeddings but also distinguishes the corresponding labels.
(2) $\min_{D}\mathbb{E}_{\boldsymbol{h},y\sim P_{\boldsymbol{H}_{aux},Y}}CE(y^{+},D(\boldsymbol{h}))+\mathbb{E}_{\boldsymbol{h},y\sim P_{\boldsymbol{H}_{train},Y}}CE(y^{-},D(\boldsymbol{h})).$
Here, $P_{\boldsymbol{H}_{aux},Y}$ and $P_{\boldsymbol{H}_{train},Y}$ represent the joint distributions of embeddings and labels from the auxiliary and training datasets, respectively. The labels $y^{+}$ and $y^{-}$ denote real and fake labels, respectively, and $\text{CE}(\cdot)$ represents the cross-entropy loss.
To align the target model’s feature embeddings more closely with the encoder’s embedding, the adversary modifies the honest training task into an adversarial task:
(3) $\min_{\boldsymbol{h}}\mathbb{E}_{\boldsymbol{h},y\sim P_{\boldsymbol{H}_{train},Y}}CE(y^{+},D(\boldsymbol{h})).$
Incorporating DAC with this adversarial task minimizes the Kullback–Leibler (KL) divergence between $P_{\boldsymbol{H}_{aux},Y}$ and $P_{\boldsymbol{H}_{train},Y}$, i.e., $\textup{KL}(P_{\boldsymbol{H}_{aux},Y}\|P_{\boldsymbol{H}_{train},Y})$. This integration ensures that the adversarial task remains indistinguishable from the honest training task. The correlation of generated gradients with labels not only reduces the detectability, but also enhances the effectiveness of the embedding distribution transfer and thus the reconstruction performance.
Comparison between discriminator and DAC: We carry out empirical studies on the distribution transfer capabilities of discriminator and the DAC, via experiments on the Credit dataset with 2 labels, and the MNIST dataset with 10 labels. We use t-distributed stochastic neighbor embedding (t-SNE) to visualize the embedding distributions, and quantify the alignment of the target model’s embeddings with those of the encoder by measuring the average cosine distance between them.
The results, illustrated in Figure 4 and 5, indicate a superior performance of the DAC in distribution transfer over discriminator. The t-SNE visualizations reveal that the embeddings’ distribution of the target model more closely resembles that of the encoder when the DAC is employed. Furthermore, a smaller average cosine distance between the embeddings of the target model and the encoder is observed when using the DAC.
(a) Encoder’s embedding distribution.
(b) Target model’s embedding distribution using discriminator. Average cosine distance is 0.6121.
(c) Target model’s embedding distribution using DAC. Average cosine distance is 0.1373.
Figure 4. t-SNE visualization on Credit dataset with 2 classes.
(a) Encoder’s embedding distribution.
(b) Target model’s embedding distribution using discriminator. Average cosine distance is 0.0412.
(c) Target model’s embedding distribution using DAC. Average cosine distance is 0.0283.
Figure 5. t-SNE visualization on MNIST dataset with 10 classes.
Malicious gradient generation using DAC: During each training round, both the adversary and other clients select a batch of training data, denoted as $\mathcal{B}_{train}$. Upon receiving the embeddings $\boldsymbol{h}_{j,p},j\in\mathcal{B}_{train}$ from passive clients, the adversary computes the malicious loss defined by:
(4) $\mathcal{L}_{M}=\frac{1}{|\mathcal{B}_{train}|}\sum_{j\in\mathcal{B}_{train}}CE\left(y_{j}^{+},D(\boldsymbol{h}_{j,p})\right).$
[…]
(5) $\mathcal{L}_{D}$ $=\frac{1}{|\mathcal{B}_{train}|}\sum_{j\in\mathcal{B}_{train}}CE(y_{j}^{-},D(\boldsymbol{h}_{j,p}))$
$+\frac{1}{|\mathcal{B}_{aux}|}\sum_{i\in\mathcal{B}_{aux}}CE(y_{i}^{+},D(\boldsymbol{h}_{i,p})).$
Minimizing $\mathcal{L}_{D}$ helps the DAC to better differentiate between real and fake embeddings associated with their labels, thereby refining its ability to guide the learning process of the target model while staying undetected. The detailed steps of malicious gradient generation are given in Algorithm 2.
Data: Auxiliary dataset $D_{aux}=\{\boldsymbol{x}_{i,a},\boldsymbol{x}_{i,p}\}_{i=1}^{M_{aux}}$ and training dataset $D_{train}=\{\boldsymbol{x}_{j,a},\boldsymbol{x}_{j,p}\}_{j=1}^{M}$.
Initialization: Encoder $f_{e}(\cdot)$, DAC $D(\cdot)$, and target model $f_{p}(\cdot)$.
Adversary Procedure:
while VFL training do
      All clients agree a batch data $\mathcal{B}_{train}$ from $D_{train}$;
      Receive and record passive clients’ embeddings $\boldsymbol{h}_{j,p},j\in\mathcal{B}_{train}$;
      Compute the loss $\mathcal{L}_{M}=\frac{1}{|\mathcal{B}_{train}|}\sum_{j\in\mathcal{B}_{train}}CE(y_{j}^{+},D(\boldsymbol{h}_{j,p}))$;
[…]
      Send $\boldsymbol{\nabla}_{p}$ to the passive clients;
      The adversary select a batch of data $\mathcal{B}_{aux}$ from $D_{aux}$;
      Compute $\boldsymbol{h}_{i,p}=f_{e}(x_{i,p}),i\in\mathcal{B}_{aux}$;
      Compute the DAC loss $\mathcal{L}_{D}$ in (5);
[…]
      $D(\cdot)\leftarrow\text{Model\_update}(D(\cdot),\boldsymbol{\nabla}_{D}$);
end while
Passive Clients Procedure:
while VFL training  do
      All clients agree a batch data $\mathcal{B}_{train}$ from $D_{train}$;
      Compute embeddings $\boldsymbol{h}_{j,p}=f_{p}(\boldsymbol{x}_{j,p}),j\in\mathcal{B}_{train}$ and send embeddings to the active client;
      Receive gradient $\boldsymbol{\nabla}_{p}$;
      $f_{p}(\cdot)\leftarrow\text{Model\_update}(f_{p}(\cdot),\boldsymbol{\nabla}_{p})$;
end while
Algorithm 2 Malicious gradient generation.
[…]
5.4. Data reconstruction
Having maliciously guided the training of the target model $f_{p}(\cdot)$ to mimic the encoder $f_{e}(\cdot)$, the adversary leverages the trained decoder $f_{d}(\cdot)$, alongside its bottom model $f_{a}(\cdot)$ to perform data reconstruction. For a target sample $\boldsymbol{x}$ (from training or test set) with adversary’s partial features $\boldsymbol{x}_{a}$ and passive clients’ features $\boldsymbol{x}_{p}$, the adversary can reconstruct the private target features $\tilde{\boldsymbol{x}}_{t}=f_{d}(f_{a}(\boldsymbol{x}_{a})\|f_{p}(\boldsymbol{x}_{p}))$.
6. Experiments
6.1. Experiment setup
Datasets and data processing: We conduct experiments across four representative datasets for VFL tasks, including two tabular datasets (i.e., Credit (Yeh, 2016) and RT_IOT2022 (S. and Nagapadma, 2024)) and two image datasets (i.e., MNIST (Deng, 2012) and CIFAR-10 (Krizhevsky et al., 2009)). Each dataset is partitioned into auxiliary, training, and test sets. For MNIST and CIFAR-10, we use the standard test sets provided. For the Credit dataset, we randomly allocate 30% of the samples to the test set, and for RT_IOT2022, 20% of the samples are selected as the test set. The auxiliary set, chosen randomly from the original training set, comprises 10% of the training set size in all main experiments, ensuring $|D_{aux}|:|D_{train}|=1:10$ and $D_{aux}\cap D_{train}=\emptyset$. Data features are evenly split between passive (target) and active clients, with each holding 50% of the features.
Models: The adversary’s encoder $f_{e}(\cdot)$ is configured identically to the passive clients’ bottom model $f_{p}(\cdot)$ across all datasets. The structure of the adversary’s bottom model $f_{a}(\cdot)$ mirrors that of $f_{p}(\cdot)$. For tabular data, we employ multi-layer perceptron (MLP) models for $f_{e}(\cdot),f_{a}(\cdot)$, $f_{p}(\cdot)$, the decoder $f_{d}(\cdot)$, and the DAC $D(\cdot)$. For image data, convolutional neural networks (CNNs) serve as the foundational architecture, with the CIFAR-10 models featuring a residual block composed of three CNN layers and a skip connection. Detailed dataset information and model structures are provided in the Appendix B.
Metrics: We measure the reconstruction performance using average MSE between the reconstructed features and the original features, denoted as Recon MSE. To analyze the effectiveness of DAC compared to other embedding distribution transfer techniques (e.g., using discriminator or the top model), we measure both the average MSE and the cosine distance between the encoder’s (or shadow model’s) embeddings and the target model’s embeddings, denoted as Emb MSE and Emb Cos, respectively. Before computing the cosine distance, all embeddings are normalized.
Environment: All experiments are conducted on a single machine equipped with four NVIDIA RTX 3090 GPUs. Each experiment is repeated five times, with the average results reported.
Table 1. Results of data reconstruction attacks across four datasets. Attack methods followed by SG or GS represent implementations under SG or GS defenses. DMAVFL and DMAVFL_sync achieve the same performance with and without defenses. Some entries are missing due to 1) some attacks do not use encoders for data reconstruction, and hence do not have embedding MSE and cosine distance; and 2) the GS defense does not work with Credit dataset.
Method Credit RT_IOT2022 MNIST CIFAR-10
Recon
MSE
Emb
MSE
Emb
Cos
Recon
MSE
Emb
MSE
Emb
Cos
Recon
MSE
Emb
MSE
Emb
Cos
Recon
MSE
Emb
MSE
Emb
Cos
GRNA 1.2566 - - 2.2955 - - 0.4209 - - 0.3201 - -
GIA 0.9054 0.2748 0.4949 1.8523 1.3960 0.4077 0.9173 0.0095 0.2535 0.2547 1.1431 0.3210
AGN 1.0934 - - 2.3671 - - 0.0929 - - 0.4549 - -
AGN-SG 1.5348 - - 3.6687 - - 0.5067 - - 0.6535 - -
AGN-GS - - - 4.0843 - - 0.5064 - - 0.4559 - -
PCAT 0.8612 0.5282 0.5486 2.2177 2.3307 0.3817 0.3694 0.0461 0.6016 0.3844 2.4672 0.5544
SDAR 0.5327 0.5451 0.3103 1.6084 2.9362 0.2281 0.0839 0.1638 0.3142 0.1162 1.5072 0.3465
FSHA 0.5152 0.3468 0.1524 1.5150 3.0500 0.4929 0.0156 0.0030 0.0378 0.0317 0.4657 0.1610
FSHA-SG 0.6034 0.5528 0.4043 1.6817 4.3649 0.5839 0.4674 0.0578 0.2016 0.5136 3.5306 0.5516
FSHA-GS - - - 1.8643 2.8935 0.3071 0.5029 0.2879 0.1760 0.1029 1.4048 0.2554
DMAVFL (GS/SG) 0.4389 0.2918 0.1492 1.2789 2.5623 0.0833 0.0152 0.0041 0.0289 0.0296 0.4348 0.1007
DMAVFL_sync (GS/SG) 0.4602 0.3010 0.1529 1.2324 2.1772 0.1025 0.0137 0.0045 0.0388 0.0305 0.4727 0.1121
6.2. Baselines and defenses
To evaluate the efficacy of our methods, we compare with three prominent data reconstruction strategies in VFL, as well as adaptations of attacks from SL. For HBC settings in VFL, we select GRNA and GIA, and for the malicious setting, we consider AGN. Additionally, we adapt three SL strategies for VFL: HBC attacks PCAT and SDAR, and the malicious attack FSHA.
GRNA: The adversary uses a generator model to create synthetic features. These synthetic features are then used to query the target model to update the generator, which minimizes the distance between the predictions made from synthetic features and those from real features.
GIA: This approach involves training a shadow model to emulate the target model’s behavior. Then the adversary feeds random noise into the trained shadow model and optimizes noise to reduce the distance between the noise embeddings and real data embeddings.
AGN: AGN discards the top model and instead uses a generator to produce synthetic data from the adversary’s features and aggregated embeddings. Concurrently, a discriminator works to differentiate between the generator’s output and genuine data samples.
PCAT: PCAT uses the top model to guide a shadow model, which simulates the behavior of the target model. A decoder is trained on the shadow model to reconstruct the target features.
SDAR: SDAR builds on PCAT by adding two discriminators that evaluate both the embeddings from the shadow model and the reconstructed features.
FSHA: In FSHA, an embedding discriminator replaces the top model. Besides, a shadow model and a decoder are trained to reconstruct target features.
We also implement SOTA defensive methods, SplitGuard (SG) and Gradient Scrutinizer (GS) on malicious attacks, to assess reconstruction performance under these defenses.
SplitGuard (SG): Analyzes gradients from fake and regular batches to compute a detection score after each round of fake batches. Following the original method, the target client calculates an average score from the last 10 records; if it falls below a threshold of 0.9, an attack is detected, and updates to the bottom model are halted.
Gradient Scrutinizer (GS): Evaluates the detection score of received gradients in each round to identify potential attacks. Due to instability in detection scores leading to frequent misjudgments, we adopt an average score strategy to improve reliability and set the threshold to 0.8 in our experiments. Similarly, when the average detection score falls below the detection threshold, an attack is detected.
In our experimental results, we append the GS or SG to the names of baseline methods to indicate their evaluation under defense mechanisms, SG or GS. We add (GS/SG) to our methods indicating the evaluation of the attack with and without defenses, GS and SG.
6.3. Data reconstruction results
Overview: As shown in Table 1, for all datasets, our proposed DMAVFL and DMAVFL_sync consistently outperform all other approaches in achieving the lowest reconstruction errors and embedding cosine distance. This underscores the effectiveness of our techniques and the advantage of embedding distribution transfer using DAC. For the RT_IOT2022 dataset, DMAVFL_sync achieves a reconstruction error of 1.2324 and cosine distance of 0.1025, significantly outperforming FSHA, which has a reconstruction error of 1.5150 and cosine distance of 0.4929.
Our methods demonstrate robustness against defensive mechanisms, maintaining consistent performance and circumventing both tested defense methods for malicious attacks. In sharp contrast, the performance of AGN and FSHA become worse when faced with the defenses, and even completely failed under some circumstances. Specifically, AGN’s reconstruction error increases dramatically from 0.0929 to 0.5067, and FSHA’s from 0.0153 to 0.5029 for the MNIST dataset. This demonstrates the stealthiness and adaptability of DMAVFL and DMAVFL_sync compared to other methods.
An important observation from our experiments is that a lower reconstruction error does not necessarily correspond to a smaller embedding distance. For example, for the MNIST dataset, while DMAVFL_sync achieves a lower reconstruction error than DMAVFL, it exhibits a slightly higher MSE and cosine embedding distance. This can be attributed to the synchronous training of all model components in DMAVFL_sync, which promotes dynamic learning and results in model diversity. Despite the higher average embedding distance, the decoder in DMAVFL_sync is more effective at reconstructing features, effectively compensating for the increased distance. Additionally, a smaller MSE embedding distance does not guarantee a more accurate reconstruction or a more similar embedding structure, as evidenced by GIA in the Credit dataset and FSHA in the MNIST dataset, which has the smallest MSE embedding distance but significantly higher cosine embedding distance and reconstruction error compared to other methods.
We do not record AGN-GS and FSHA-GS in the Credit dataset. It is because GS doesn’t work in the Credit dataset, which can also be observed in Figure 6(h), caused by the limited data features (23 features) and imbalanced label distribution (77% label 0 and 23% label 1).
Visualization of the reconstruction: Table 2 visualizes the reconstruction performance of the proposed DMAVFL, alongside the malicious attack baselines, AGN and FSHA, for the CIFAR-10 dataset. The results clearly demonstrate that DMAVFL produces more accurate and distinct reconstructions compared to the baselines. While AGN and FSHA yield coarser results, DMAVFL maintains high-quality reconstruction even under defensive conditions. In contrast, both FSHA and AGN struggle to reconstruct informative pixels effectively when faced with defense mechanisms.
Table 2. Visualization of reconstructed features on the CIFAR-10 dataset with and without the defense. Except for the original image, the right side of each image is the reconstructed image. The last column is the average reconstruction error.
SG and GS defense results: Figure 6 illustrates the detection scores for DMAVFL and DMAVFL_sync, which consistently exceed the detection thresholds, aligning indistinguishably with honest training scenarios. Although in the CIFAR-10 dataset, DMAVFL’s detection score is lower than the threshold at the beginning, its average score remains above the threshold, ensuring stealthiness against target clients. In contrast, malicious attack baselines, AGN and FSHA, consistently maintain scores below the detection thresholds of SG and GS in the MNIST and CIFAR-10 datasets, rendering them detectable.
In the RT_IOT2022 and Credit datasets under defense SG, AGN sporadically surpasses the threshold, yet its average detection score falls below the threshold, leading passive clients to detect an attack promptly. FSHA’s SG scores remain below the threshold throughout, making it easily detectable. Under GS, FSHA exceeds the threshold at the beginning but is ultimately detected in the RT_IOT2022 dataset. In the Credit dataset, as previously noted, GS is ineffective due to imbalanced labels and limited features.
(a) SG scores in MNIST.
(b) SG scores in CIFAR-10.
(c) SG scores in RT_IOT2022.
(d) SG scores in Credit.
(e) GS scores in MNIST.
(f) GS scores in CIFAR-10.
(g) GS scores in RT_IOT2022.
(h) GS scores in Credit.
Figure 6. Detection scores of SG and GS on malicious attacks and the honest training across four datasets.
7. Ablation Study
This section explores the robustness of DMAVFL and DMAVFL_sync (abbreviated as sync) by evaluating their performance under various parameter settings. These settings include changes in the adversary’s feature size, use of heterogeneous encoder models, variations in auxiliary dataset size, and differences in the number of target clients. Except for the variables specifically altered for this study, the model setups and implementation environment remain consistent with those described in the main experiments.
7.1. Effect of adversary’s feature size
To explore how the size of the feature set held by the adversary affects the reconstruction performance of DMAVFL and sync, we conduct experiments targeting the reconstruction of 30% of the features while varying the proportion of features held by the adversary from 0% to 70%. At 0% feature possession, the scenario reduces to split learning where no features or bottom model are held by the adversary. These experiments are performed on the Credit and RT_IOT2022 (abbreviated as IOT) datasets, with the results summarized in Table 3.
Table 3. Reconstruction error under various adversary’s feature portions.
Feature portion 0% 10% 20% 50% 70%
Credit DMAVFL 0.8315 0.7839 0.7489 0.6987 0.5189
sync 0.8718 0.7537 0.7154 0.6940 0.5732
IOT DMAVFL 1.7423 1.7014 1.6939 1.6382 1.5456
sync 1.7848 1.6721 1.6103 1.6056 1.2811
Consistent with our expectations, increasing the proportion of features held by the adversary leads to improved reconstruction accuracy. As illustrated in Table 3, the reconstruction error in the Credit dataset using DMAVFL decreases from 0.8315 to 0.5189 as the feature possession increases from 0% to 70%. This trend is observed for both datasets and methods.
7.2. Effect of encoder’s complexity
In our main experiments, the encoder mirrors the target model’s structure. However, when the adversary lacks precise knowledge about the target model’s configuration, discrepancies between the encoder and the target model’s structures can affect the transfer of embedding distributions and, consequently, the overall reconstruction performance. To evaluate the impact of varying encoder complexities, we conduct experiments on the CIFAR-10 dataset, where the target model comprises 3 residual blocks. We test the following encoder configurations to assess their effect on performance. MLP Model: Comprises 3 linear layers, with input and output dimensions matching those of the target model; 1-Res Model: Utilizes a single residual block. 5-Res Model: Extends to 5 residual blocks, increasing complexity. We test the reconstruction error and measure both the embeddings MSE and cosine distances for each encoder configuration.
Table 4. Reconstruction error and average embeddings distance using different encoder’s model.
Encoder Metric DMAVFL sync
MLP Recon MSE 0.2545 0.2362
Emb MSE dis 7.7767 3.9726
Emb Cos dis 0.6337 0.6081
1-Res Recon MSE 0.0561 0.0434
Emb MSE dis 0.3175 0.3059
Emb Cos dis 0.1632 0.1450
3-Res (Same) Recon MSE 0.0295 0.0305
Emb MSE dis 0.4347 0.4727
Emb Cos dis 0.1006 0.1121
5-Res Recon MSE 0.0313 0.0366
Emb MSE dis 0.8771 0.7071
Emb Cos dis 0.1081 0.1112
Table 4 clearly illustrates that the adversary achieves the lowest reconstruction error and the smallest embedding cosine distance when using an encoder the same as the target model, as observed for both DMAVFL and sync. This optimal alignment underscores the importance of model congruence in enhancing the effectiveness of embedding distribution transfer and data reconstruction. Conversely, when the encoder structure deviates from that of the target model, specifically when switching from a CNN to an MLP, the reconstruction performance significantly deteriorates, with reconstruction error increasing by an order of magnitude (from 0.0295 to 0.2545). This dramatic drop highlights the challenges of using less compatible models for data reconstruction attack. Moreover, the results indicate that a complex encoder, featuring a higher number of residual blocks, outperforms a simpler encoder configuration. The complex encoder not only enhances the decoder’s ability to reconstruct the data accurately but also learns a more informative embedding. This suggests that increasing the encoder’s complexity, when the target model is unknown, can lead to better performance in both embedding distribution transfer and data reconstruction.
7.3. Effect of auxiliary dataset
In the main experiments, we set the size of the adversary’s auxiliary dataset, $|D_{aux}|$, relative to the training set size, $|D_{train}|$, at a ratio of $1:10$. To explore the impact of varying auxiliary dataset sizes, we adjust the ratio of $|D_{aux}|:|D_{train}|$ from $0:10$ to $1:1$ in the MNIST dataset. Specifically, at a ratio of 0:10, the auxiliary dataset comprises the FashionMNIST (Xiao et al., 2017) training dataset instead of a subset of MNIST, introducing a distinct distribution from the training set. This setup aims to test the adaptability of the reconstruction process to changes in the auxiliary dataset’s distribution.
Table 5. Reconstruction error for different $|D_{aux}|$ relative to $|D_{train}|$.
$|D_{aux}|:|D_{train}|$ 0:10 1:20 1:10 1:5 1:2 1:1
DMAVFL 0.1076 0.0315 0.0152 0.0186 0.0188 0.0172
sync 0.2311 0.0240 0.0137 0.0166 0.0178 0.0189
From Table 5 we can see that significant data leakage can occur, even in the absence of auxiliary data with the same distribution of the training set (i.e., with out-of-distribution auxiliary data). This observation is highlighted by a reconstruction error of 0.1076$\sim$0.2311, when the adversary uses FashionMNIST data to reconstruct images from MNIST. We also observe that for both DMAVFL and sync, the reconstruction error decreases quickly as $|D_{aux}|$ becomes relatively larger. However, after the ratio of $|D_{aux}|:|D_{train}|$ reaches 1:10, increasing auxiliary dataset size results in minor improvements in reconstruction error. This suggests that there is a threshold beyond which additional auxiliary data does not substantially improve the model’s ability to reconstruct target features, pointing to a diminishing return on the utility of expanding the auxiliary set.
7.4. Effect of the number of target clients
Our proposed methods, DMAVFL and sync, are designed to handle scenarios involving multiple target clients. To assess the impact of varying the number of target clients on data reconstruction performance, we conduct experiments on the IOT dataset with the number of target clients ranging from 1 to 5. Consistent with the setup in the main experiments, 50% of the features are equally distributed among the target clients.
Table 6. Reconstruction errors for different numbers of target clients.
# of target clients 1 2 3 4 5
DMAVFL 1.2789 1.5801 1.5986 1.6718 1.9820
sync 1.2324 1.6447 1.7597 1.8660 2.2635
The results presented in Table 6 indicate that an increase in the number of target clients, each sharing a smaller portion of the features, leads to diminished reconstruction performance. Specifically, the reconstruction error increases from 1.2789 with a single target client to 1.9828 when there are five target clients. This degradation in performance can be attributed to the distribution of features among more clients, each employing a separate bottom model to learn their respective features. This setup tends to cause the loss of crucial feature correlation information during the learning process. Moreover, as the encoder is trained on the complete set of features and is hence able to capture more feature correlations, the increased disparity between the encoder and the target models’ understanding of the feature correlation further exacerbates the challenge of accurate data reconstruction.
8. Evaluations on other defenses
While DMAVFL is designed and demonstrated to penetrate detection-based defenses of SplitGuard and Gradient Scrutinizer, we also evaluate its effectiveness against general defensive methods for privacy protection in VFL and SL. Note that unlike SplitGuard and Gradient Scrutinizer, these methods essentially introduce perturbations to the communicated embeddings or gradients, severely degrading the performance of honest training. Specifically, we consider three such defenses: Nopeek (Vepakomma et al., 2020), obfuscation with random noise (He et al., 2020), and differential privacy (DP) (Dwork, 2006; Abadi et al., 2016; Wei et al., 2020).
Nopeek Defense. Nopeek is a method for minimizing distance correlation, widely implemented in both VFL and SL to prevent privacy leakage. It operates by minimizing the correlation between embeddings and the original features to reduce information leakage. During VFL training, passive clients incorporate the Nopeek loss into their training loss, defined as:
(6) $\mathcal{L}_{N}$ $=\frac{1}{|\mathcal{B}_{train}|}\sum_{j\in\mathcal{B}_{train}}(\alpha\cdot DCOR(\boldsymbol{x}_{j,n},f_{n}(\boldsymbol{x}_{j,n}))$
$+(1-\alpha)\cdot TASK(y_{j},f_{n}(\boldsymbol{x}_{j,n})),$
where $DCOR$ represents the distance correlation metric, $TASK$ denotes the VFL training loss, and $\alpha$ is a hyper-parameter balancing training and defense.
We challenge Nopeek by deploying DMAVFL and sync against it with varying $\alpha$ values from 0.1 to 1.0 on the MNIST dataset. The results in Table 7 show that when $\alpha<0.9$, Nopeek fails to thwart our attacks. As $\alpha$ approaches 0.9 and beyond, Nopeek begins to impede the attack effectiveness. Notably, even at $\alpha=0.95$, DMAVFL still achieves a small reconstruction error of 0.0280. Setting $\alpha=1.0$ effectively defends the attack. However, this completely replaces the original VFL task with minimizing the $DCOR$ loss, which results in an accuracy of merely 69.03%.
Table 7. Reconstruction error against Nopeek.
$\alpha$ 0 0.2 0.4 0.6 0.8 0.9 0.95 1.0
DMAVFL 0.0152 0.0151 0.0122 0.0138 0.0116 0.0171 0.0280 0.4759
sync 0.0137 0.0143 0.0122 0.0113 0.0135 0.0156 0.0403 0.4462
Honest Acc(%) 99.11 97.82 98.28 98.04 97.90 95.03 94.72 69.03
Obfuscation with random noise. Perturbation methods, such as adding noise to transmitted embeddings, are commonly employed in ML to safeguard privacy. Following the approach outlined in (He et al., 2020), we introduce noise with standard deviations $\sigma$ ranging from 0.2 to 0.8 into the embeddings sent by target clients. This defense is tested against DMAVFL and sync on the MNIST dataset, and the results are reported in Table 8. The results suggest that higher noise levels more effectively impede the attacks. This defense is more effective for sync, as its DAC learns a dynamic distribution of the encoding embedding during training, and is more susceptible to fluctuations caused by noise. Nonetheless, our attacks manage to reconstruct informative features even under significant noise, for instance, DMAVFL achieving a reconstruction error of 0.0472 at $\sigma=0.7$. It is also important to note that adding larger noise adversely affects the honest training accuracy.
Table 8. Reconstruction error against obfuscation.
$\sigma$ 0 0.1 0.3 0.5 0.7
DMAVFL 0.0152 0.0214 0.0252 0.0351 0.0472
sync 0.0137 0.0161 0.0301 0.0497 0.0752
Honest Acc(%) 99.11 98.34 97.65 94.84 94.00
Differential privacy (DP). DP provides a mathematical framework for quantifying privacy guarantees for ML models. In our attacks, as the gradients returned from the active client may be maliciously generated, we propose to apply $\epsilon$-DP mechanism, which involves adding Laplacian noise to received gradients, to disrupt the malicious training. We test varying levels of privacy by adjusting the parameter $\epsilon$ from 0.1 to 10, where a smaller $\epsilon$ implies stronger privacy due to higher noise levels. The average reconstruction errors of DMAVFL and sync on MNIST are shown in Table 9, indicating that DP is most effective at an $\epsilon$ of 0.1, successfully impeding our proposed attacks. However, as $\epsilon$ increases, its defensive effectiveness decreases. At $\epsilon=10$, the noise assists the attack, potentially by making the model updates more robust, thereby facilitating the embedding distribution transfer. It’s important to note that while DP helps to protect privacy, it also hurts the accuracy of honest training.
Table 9. Reconstruction error against DP. $\epsilon=+\infty$ corresponds to the scenario without DP.
$\epsilon$ 0.1 0.5 1 2 5 10 $\infty$
DMAVFL 0.3134 0.0773 0.0443 0.0210 0.0210 0.0122 0.0152
sync 0.3178 0.1253 0.0507 0.0468 0.0172 0.0110 0.0137
Honest Acc(%) 87.75 90.00 94.97 95.90 96.72 97.36 99.11
9. Conclusion
We introduce DMAVFL, a novel defense-penetrating malicious data reconstruction attack for vertical federated learning (VFL). DMAVFL innovatively incorporates training of a discriminator with auxiliary classifier (DAC) to integrate label information into the attack process, which not only makes the malicious training indistinguishable from honest training, but also substantially improves data reconstruction performance. Through visualization and quantification, we demonstrated the effectiveness of DAC compared to traditional discriminators, showcasing its superior capability in embedding transfer. Through extensive experiments, we empirically demonstrated that DMAVFL successfully evades various defensive methods and outperforms all existing data reconstruction attacks. While traditional defense methods based on embedding/model perturbations are shown to effectively defend DMAVFL, it is at the cost of serious degradation on the honest VFL task’s performance. This calls for future research to develop effective defenses against malicious data reconstruction, with minimal impact on the honest task.
Appendix A DMAVFL_sync algorithm
We detail the steps of DMAVFL_sync in Algorithm 3.
Data: Auxiliary dataset $D_{aux}=\{\boldsymbol{x}_{i,a},\boldsymbol{x}_{i,p}\}_{i=1}^{M_{aux}}$ and training dataset $D_{train}=\{\boldsymbol{x}_{j,a},\boldsymbol{x}_{j,p}\}_{j=1}^{M}$.
Initialization: Encoder $f_{e}(\cdot)$, DAC $D(\cdot)$, adversary’s bottom model $f_{a}(\cdot)$, and target model $f_{p}(\cdot)$.
Adversary Procedure:
while VFL training do
      The adversary select a batch $\mathcal{B}_{aux}$ from $D_{aux}$;
      Compute $\boldsymbol{h}_{i,a}=f_{a}(\boldsymbol{x}_{i,a}),i\in\mathcal{B}_{aux}$, and $\boldsymbol{h}_{i,p}=f_{e}(\boldsymbol{x}_{i,p}),i\in\mathcal{B}_{aux}$;
      Compute $\tilde{\boldsymbol{x}}_{i,t}=f_{d}([\boldsymbol{h}_{i,a}\|\boldsymbol{h}_{i,p}]),i\in\mathcal{B}_{aux}$;
      $\mathcal{L}_{R}\leftarrow\frac{1}{|\mathcal{B}_{aux}|}\sum_{i\in\mathcal{B}_{aux}}MSE(\tilde{\boldsymbol{x}}_{i,t},\boldsymbol{x}_{i,t})$;
      $\boldsymbol{\nabla}_{e}\leftarrow$ Gradient($\mathcal{L}_{R},f_{e}(\cdot)$);
      $\boldsymbol{\nabla}_{a}\leftarrow$ Gradient($\mathcal{L}_{R},f_{a}(\cdot)$);
      $\boldsymbol{\nabla}_{d}\leftarrow$ Gradient($\mathcal{L}_{R},f_{d}(\cdot)$);
      $f_{e}(\cdot)\leftarrow\text{Model\_update}(f_{e}(\cdot),\boldsymbol{\nabla}_{e}$);
      $f_{a}(\cdot)\leftarrow\text{Model\_update}(f_{a}(\cdot),\boldsymbol{\nabla}_{a}$);
      $f_{d}(\cdot)\leftarrow\text{Model\_update}(f_{d}(\cdot),\boldsymbol{\nabla}_{d}$);
      Compute $\boldsymbol{h}_{i,p}=f_{e}(x_{i,p}),i\in\mathcal{B}_{aux}$;
      All clients agree a batch data $\mathcal{B}_{train}$ from $D_{train}$;
      Receive and record passive clients’ embeddings $\boldsymbol{h}_{j,p},j\in\mathcal{B}_{train}$;
      Compute the loss $\mathcal{L}_{M}=\frac{1}{|\mathcal{B}_{train}|}\sum_{j\in\mathcal{B}_{train}}CE(y_{j}^{+},D(\boldsymbol{h}_{j,p}))$;
[…]
      Send $\boldsymbol{\nabla}_{p}$ to the passive clients;
      Compute the DAC loss $\mathcal{L}_{D}$ in (5);
[…]
      $D(\cdot)\leftarrow\text{Model\_update}(D(\cdot),\boldsymbol{\nabla}_{D}$);
end while
Passive Clients Procedure:
while VFL training  do
      All clients agree a batch data $\mathcal{B}_{train}$ from $D_{train}$;
      Compute embeddings $\boldsymbol{h}_{j,p}=f_{p}(\boldsymbol{x}_{j,p}),j\in\mathcal{B}_{train}$ and send embeddings to the active client;
      Receive gradient $\boldsymbol{\nabla}_{p}$;
      $f_{p}(\cdot)\leftarrow\text{Model\_update}(f_{p}(\cdot),\boldsymbol{\nabla}_{p})$;
end while
Algorithm 3 DMAVFL_sync.
Appendix B Datasets and model descriptions
We illustrate all datasets and models information used in the experiments in the following Table 10, where MLP is composed of a linear and an activation, e.g., ReLU layer, CNN layer is composed a convolution process and a linear layer, and Residual block is constructed by 3 CNN layers and a skip connection.
Table 10. Dataset characteristics and models structure
Dataset Credit RT_IOT2022 MNIST CIFAR-10
Feature size
$d_{a},d_{p}$
11, 12 41, 42 392, 392 512, 512
The size of
training set
19,091 89,920 54,546 45,455
The size of
auxiliary set
1,909 8,573 5,454 4,545
The size of
test set
9,000 24,624 10,000 10,000
$f_{e}(\cdot),f_{a}(\cdot),$ and
$f_{p}(\cdot)$ structure
MLP layer
2
MLP layer
2
CNN layer
2
Residual
block *4
$f_{d}(\cdot)$ structure
MLP layer
3
MLP layer
3
CNN layer
3
CNN layer
3
$D(\cdot)$ structure
MLP layer
3
MLP layer
3
CNN layer
2
CNN layer
4